You are the manager of a firm that receive revenue of Rs.30,000 per year from product X and Rs. 70,000 per year from product Y. The own price elasticity of demand for product X is -2.5 and the cross price elasticity of demand between product Y and X is 1.1. How much will you firm’s total revenue (revenues from both products) change if you increase the price of good X by 1 present?


1
Expert's answer
2021-07-22T10:54:44-0400

The change in firm’s total revenue can be calculated as follows:

Change in total revenue is given by:

(TRX(1+price elasticity)+TRY(cross price elasticity))*percentage change in price

"=(30,000(1+(-2.5))+70,000(1.1))(0.01)\\=(-45,000+77,000)(0.01)\\=32000\times0.01\\=320"